OpenStudy (anonymous):

Solve P = 1/6(r + 5t) for r. You must show all work.

OpenStudy (diyadiya):

\[P = \frac{1}{6}(r+5t)~~~~~or~~~~~ P= \frac{1}{6(r+5t)}\]

OpenStudy (anonymous):

is this the answer?

OpenStudy (diyadiya):

loli'm asking which is the question?

OpenStudy (anonymous):

1st one

OpenStudy (diyadiya):

Multiply both sides by 6

OpenStudy (diyadiya):

what do you get now?

jhonyy9 (jhonyy9):

P= r/6 + 5t/6 P -5t/6 =r/6 (6P -5t)/6 =r/6 6P -5t=r

OpenStudy (anonymous):

P = 1 / 6(r + 5t) 6P(r + 5t) = 1 r + 5t = 1 / 6P r = (1 / 6P) - 5t

OpenStudy (anonymous):

is that correct?

OpenStudy (diyadiya):

If your question is \[P= \frac{1}{6(r+5t)}\]Then you're right! if your question is \[P = \frac{1}{6}(r+5t)\]then you're wrong

jhonyy9 (jhonyy9):

so how you have wrote so than r+5t is the numerator and just 6 is the denominator yes ?

OpenStudy (anonymous):

Well I'm Wrong . @jhonyy9 yours is correct.

jhonyy9 (jhonyy9):

ok so than there is how you can solve it right ok ?

OpenStudy (anonymous):

Huh?

jhonyy9 (jhonyy9):

6P-5t=r

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

@Diyadiya is this the answer?

OpenStudy (anonymous):

P= r/6 + 5t/6 P -5t/6 =r/6 (6P -5t)/6 =r/6 6P -5t=r

OpenStudy (diyadiya):

Yes! you're right!

OpenStudy (anonymous):

Ok, Thanks!